Let the joint probability density function of X and Y be bivariate normal. For what values of a is the variance of X + Y minimum ?
Answer #1

Var(\alpha X+Y)=Var(\alpha X)+Var(Y) + 2cov(\alpha X,Y)

\Rightarrow Var(\alpha X+Y)=\alpha^2Var(X)+Var(Y) + 2\alpha cov(X,Y)

\Rightarrow Var(\alpha X+Y)=\alpha^2\sigma_X^2+ \sigma_Y^2+2\alpha *\sigma_X*\sigma_Y

To find \alpha such that Var(\alpha X+Y) is minimum.

Now,

Var(\alpha X+Y)\geq 0

\Rightarrow \alpha^2\sigma_X^2+ \sigma_Y^2+ 2\alpha *\sigma_X*\sigma_Y \geq 0

\Rightarrow (\alpha \sigma_X+\sigma_Y)^2 \geq 0

\Rightarrow (\alpha \sigma_X+\sigma_Y) \geq 0

\Rightarrow \alpha \geq - \frac{\sigma_Y}{\sigma_X}